OpenStudy (anonymous):

If two sides of a triangle measure 12 and 7, which cannot be the perimeter of the triangle? A) 29 B) 37 C) 34 D) 38

OpenStudy (anonymous):

D. 38.

OpenStudy (anonymous):

5 < x < 19 Because 12+7 = 19 and 12-7 = 5

OpenStudy (anonymous):

A x = 29-19 = 10 B x = 34-19 = 15 C x = 37-19 = 18 D x = 38-19 = 19

OpenStudy (anonymous):

Thank you sooo much :3